# Locally-APN Binomials with Low Boomerang Uniformity in Odd Characteristic

### v1

@@ Abstract
Recently, several studies have shown that when $q\equiv 3\pmod{4}$, the function $F_{r}(x)=x^{r}+x^{r+\frac{q-1}{2}}$ defined over $\mathbb{F}_{q}$ is locally-APN and has boomerang uniformity at most $2$. In this paper, we extend these results by showing that if there is at most one $x\in\mathbb{F}_{q}$ with $\chi(x)=\chi(x+1)=1$ satisfying $(x+1)^{r}-x^{r}=b$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $F_{r}$ is locally-APN with boomerang uniformity at most $2$. Moreover, we study the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$, and the boomerang spectrum of $F_{2}$ when $p=3$.
Keywords. Binomials, Differential Uniformity, Differential Spectrum, Boomerang Uniformity, Boomerang Spectrum, Locally APN Functions.
Mathematics Subject Classification(2020) 94A60, 06E30
@@ 1 Introduction
where $\chi(x)=x^{\frac{q-1}{2}}$ is the quadratic character over $\mathbb{F}_{q}$. Differential uniformity, introduced by Nyberg[16], quantifies the resistance of an S-box in a block cipher to differential cryptanalysis. The definition of the differential uniformity is given below.
@@
More recently, several studies have focused on the non-APN case of $F_{q-2,u}$. For instance, Xia et al. [18] investigated the differential uniformity of $F_{q-2,u}$ and determined its differential spectrum when $\chi(u+1)=\chi(u-1)$ in characteristic $3$. Their results were later extended in [19] to the general case where $q\equiv 3\pmod{4}$. Ren et al. [17] studied the differential spectrum of $F_{q-2,u}$ in the complementary case $\chi(u+1)\neq\chi(u-1)$, while Lyn et al. [13] independently showed that $F_{q-2,u}$ is a differentially $4$-uniform permutation under the same condition.
Budaghyan and Pal [3] conjectured, based on experimental evidence, that there might exist an infinite subclass of APN functions among $F_{2,u}$. However, this conjecture was disproved in [14], where the authors conducted a detailed study of the differential uniformity of $F_{2,u}$. The same conjecture was independently disproved in [1] by means of function field theory. Koo and Kwon [10] further investigated the differential uniformity of $F_{r,u}$ for $r=\frac{q+1}{4}$. The differential properties of $F_{r,u}$ are summarized in Table 1.
@@
q2q-2 $u\neq\pm 1$ $p=3$, $n>3$ : odd $\chi(1+u)=\chi(u-1)=\chi(u)$ 2 (APN) O [18]
q2q-2 $u\neq\pm 1$ $p=3$, $n>3$ : odd $\chi(1+u)=\chi(u-1)\neq\chi(u)$ 33 O [18]
q2q-2 $u\neq\pm 1$ $p=3$, $n>3$ : odd $\chi(1+u)\neq\chi(u-1)$ $\delta_{F_{r,u}}\leq 4$ O [17]
q2q-2 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $\delta_{F_{r,u}}\leq 5$ X [3]
22 $\pm 1$ $q\equiv 3\pmod{4}$ $\frac{q+1}{4}$ (locally-APN) O [14]
22 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $3\leq\delta_{F_{r,u}}\leq 5$ O [14]
$\frac{q+1}{4}$ $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $\delta_{F_{r,u}}\leq 5$ X [10]
@@ 22 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $3\leq\delta_{F_{r,u}}\leq 5$ O [14]
$\frac{3^{k}+1}{2}$ $\pm 1$ $p=3$, $n$ : odd $\frac{q+1}{4}$ (locally-APN) X This paper
@@ 33 $\pm 1$ $q\equiv 11\pmod{12}$ $\frac{q+1}{4}$ (locally-APN) O This paper
Recently, several studies have investigated the differential and boomerang spectra of $F_{r}$ for various choices of $r$. These works consistently show that $F_{r}$ is locally-APN and has boomerang uniformity at most $2$ in all known cases. Lyn et al. [13] determined the differential and boomerang spectra of $F_{q-2}$, showing that it is locally-PN with boomerang uniformity 0 when $p=3$, and locally-APN with boomerang uniformity at most $1$ when $p\neq 3$. Mesnager and Wu [14] proved that $F_{2}$ is locally-APN with boomerang uniformity at most $2$, and further investigated its differential spectrum. They also proved that the boomerang uniformity of $F_{2}$ equals $2$, when $q\geq 9613^{2}$. However, we will show in Section 4.2 that the boomerang uniformity of $F_{2}$ equals $1$ when $p=3$. Yan and Ren [20] independently analyzed the differential spectrum of $F_{2}$ using a different approach. Koo and Kwon [10] studied $F_{r}$ for $r=\frac{q+1}{4}$, showing that $F_{r}$ is locally-PN with boomerang uniformity 0 when $q\equiv 3\pmod{8}$, and locally-APN with boomerang uniformity $2$ when $q\equiv 7\pmod{8}$. They also provided complete characterizations of the differential and boomerang spectra of $F_{r}$ in these cases. The boomerang properties of $F_{r}$ are summarized in Table 2.
@@ 22 $q\equiv 3\pmod{4}$, large $q$ 2 X [14]
$\frac{3^{k}+1}{2}$ p=3,np=3,\;n : odd $\leq 2$ X This paper
33 $q=11\;(\bmod 12)$ $\leq 2$ X This paper
$\frac{2q-1}{3}$ $q=11\;(\bmod 12)$ $\leq 2$ X This paper
@@ 22 $p=3,\;n\geq 3$ : odd 1 O This paper
In this paper, we study the differential and boomerang properties of $F_{r}$ when $q\equiv 3\pmod{4}$. In particular, we show that if the equation $(x+1)^{r}-x^{r}=b$ has at most one solution satisfying $\chi(x)=\chi(x+1)=1$ and $\gcd(r,q-1)\mid 2$, then $\delta_{F_{r}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$. This implies that $F_{r}$ is locally-APN. Furthermore, we also prove that its boomerang uniformity is at most $2$. The above condition is more likely to hold when the differential uniformity of $x^{r}$ is small, and it certainly holds when $x^{r}$ is PN. In this paper, we show that this is indeed the case for the exponents $r$ listed in Table 3 in Section 2. In addition, we further investigate the differential spectrum for $r=3$ and $r=\frac{2q-1}{3}$, and the boomerang spectrum of $F_{2}$ when $p=3$. By Lemma 2.8, the latter result can be extended to the case k=n1k=n-1 for $r=\frac{3^{k}+1}{2}$.
@@ 2 Preliminaries
When $F(x)=x^{r}$ is a power function, multiplying $\frac{1}{a^{r}}$ on the both sides of $b=F(x+a)-F(x)=(x+a)^{d}-x^{d}$, we have
@@
in this case. Notion of locally-APN functions was introduced in [2], for the even characteristic case. In [9], Hu et al generalized it for odd characteristic cases, as follows.
@@
$\delta_{F}(a,b)=\delta_{F}(1,g_{a}(b))\text{ for all }b\in\mathbb{F}_{p^{n}},$ (4)
@@
Let $F$ be a function on $\mathbb{F}_{q}$ satisfying (4) for every $a\in\mathbb{F}_{p^{n}}^{*}$. Then, $F$ is called locally-APN if $\delta_{F}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}\setminus\mathbb{F}_{p}$.
@@
[6] If $p^{n}\equiv 1\pmod{4}$, then $S_{00}=\frac{p^{n}-5}{4}$ and $S_{01}=\#S_{10}=\#S_{11}=\frac{p^{n}-1}{4}$. If $p^{n}\equiv 3\pmod{4}$, then $S_{00}=\#S_{10}=\#S_{11}=\frac{p^{n}-3}{4}$ and $S_{01}=\frac{p^{n}+1}{4}$.
@@
(Theorem 5.48 of [12]) Let $f(x)=a_{2}x^{2}+a_{1}x+a_{0}\in\mathbb{F}_{p^{n}}[x]$ with $p$ odd and $a_{2}\neq 0$. Put $d=a_{1}^{2}-4a_{0}a_{2}$. Then,
$\sum_{x\in\mathbb{F}_{p^{n}}}\chi(f(x))=\begin{cases}-\chi(a_{2})&\text{ if }d\neq 0,\\ (p^{n}-1)\chi(a_{2})&\text{ if }d=0.\end{cases}$
@@
(Theorem 5.41 of [12]) Let $\eta(\cdot)$ be a multiplicative character of $\mathbb{F}_{q}$ of order $m>1$ and let $f\in\mathbb{F}_{q}[x]$ be a monic polynomial of positive degree that is not $m$-th power of a polynomial. Let $d$ be the number of distinct roots of $f$ in its splitting field over $\mathbb{F}_{q}$. Then for every $a\in\mathbb{F}_{p^{n}}$ we have
@@
In this paper, we show that for $r$ in Table 3 $F_{r}$ is locally-APN with differential uniformity $\frac{q+1}{4}$ with boomerang uniformity at most $2$. These exponents $r$ form a subclass of those satisfying $\gcd(r,q-1)=2$ for which the monomial $x^{r}$ has low differential uniformity. Note that the differential properties summarized in Table 3 are stated for the monomial $x^{r}$, but in some cases the equivalent form $x^{r+\frac{q-1}{2}}=x^{r}\chi(x)$ is actually PN or APN. Since $F_{r+\frac{q-1}{2}}(x)=x^{r+\frac{q-1}{2}}(1+\chi(x))=x^{r}(\chi(x)+1)=F_{r}(x)$ for all $x\in\mathbb{F}_{q}$, the two forms are identical as functions.
@@ 33 $q\equiv 11\pmod{12}$ APN -
Moreover, when $r=\frac{3^{k}+1}{2}$ (the Coulter–Matthews exponent), $x^{r}$ is PN only for odd $k$ [5]. However, according to the linear equivalence established in Lemma 2.8, our results hold for all $k$. Two functions $F$ and $F^{\prime}$ defined on $\mathbb{F}_{q}$ are said to be linearly(affine) equivalent if there exist linear(affine) permutations $L_{1}$ and $L_{2}$ over $\mathbb{F}_{q}$ such that $F^{\prime}=L_{1}\circ F\circ L_{2}$. It is well known that the differential and boomerang spectra are invariant under affine equivalence.
@@
Let $p=3$ and $n$ be odd. Then $F_{\frac{3^{k}+1}{2}}$ is linearly equivalent to $F_{\frac{3^{n-k}+1}{2}}$.
@@
Let $L(x)=x^{3^{n-k}}$. Then, for nonzero $x\in\mathbb{F}_{q}$,
$(F_{\frac{3^{k}+1}{2}}\circ L)(x)$ $=\left(x^{3^{n-k}}\right)^{\frac{3^{k}+1}{2}}(1+\chi(x^{3^{n-k}}))=x^{\frac{3^{n}+3^{n-k}}{2}}(1+\chi(x))=x^{\frac{3^{n-k}+1}{2}}\cdot x^{\frac{3^{n}-1}{2}}(1+\chi(x))$
$=x^{\frac{3^{n-k}+1}{2}}(\chi(x)+1)=F_{\frac{3^{n-k}+1}{2}}(x).$
Clearly, we can easily check that $(F_{\frac{3^{k}+1}{2}}\circ L)(0)=F_{\frac{3^{n-k}+1}{2}}(0)=0$. ∎
@@ 3 Differential Properties
In this section, we study the differential properties of $F_{r}$. We desire to count the number of solutions of
@@
Denote $D_{ij}(b)$ be the number of solutions of (5) in $S_{ij}$ where $i,j\in\{0,1\}$. Then the following is almost clear.
$\delta_{F_{r}}(1,b)=\begin{cases}\sum D_{ij}(b)&\text{ if }b\neq 0,2,\\ \sum D_{ij}(b)+1&\text{ if }b=0\text{ or }2.\end{cases}$

### v2

@@ Abstract
Recently, several studies have shown that when $q\equiv 3\pmod{4}$, for certain choices of $r$, the function $F_{r}(x)=x^{r}+x^{r+\frac{q-1}{2}}$ defined over $\mathbb{F}_{q}$ is locally-APN and has boomerang uniformity at most $2$. In this paper, we extend these results by showing that if there is at most one $x\in\mathbb{F}_{q}$ with $\chi(x)=\chi(x+1)=1$ satisfying $(x+1)^{r}-x^{r}=b$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $F_{r}$ is locally-APN with boomerang uniformity at most $2$. Moreover, we study the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$, and the boomerang spectrum of $F_{2}$ when $p=3$.
Keywords. Binomials, Differential Uniformity, Differential Spectrum, Boomerang Uniformity, Boomerang Spectrum, Locally-APN Functions.
Mathematics Subject Classification(2020) : 94A60, 06E30
@@ 1 Introduction
where $\chi(x)=x^{\frac{q-1}{2}}$ is the quadratic character over $\mathbb{F}_{q}$, and $u\in\mathbb{F}_{q}^{*}$. Differential uniformity, introduced by Nyberg[16], quantifies the resistance of an S-box in a block cipher to differential cryptanalysis. The definition of the differential uniformity is given below.
@@
More recently, several studies have focused on the non-APN case of $F_{q-2,u}$. For instance, Xia et al. [19] investigated the differential uniformity of $F_{q-2,u}$ and determined its differential spectrum when $\chi(u+1)=\chi(u-1)$ in characteristic $3$. Their results were later extended in [20] to the general case where $q\equiv 3\pmod{4}$. Ren et al. [17] studied the differential spectrum of $F_{q-2,u}$ in the complementary case $\chi(u+1)\neq\chi(u-1)$, while Lyu et al. [13] independently showed that $F_{q-2,u}$ is a differentially $4$-uniform permutation under the same condition.
Budaghyan and Pal [3] conjectured, based on experimental evidence, that there might exist an infinite subclass of APN functions among $F_{2,u}$. However, this conjecture was disproved in [14], where the authors conducted a detailed study of the differential uniformity of $F_{2,u}$. The same conjecture was independently disproved in [1] by means of function field theory. Koo and Kwon [10] further investigated the differential uniformity of $F_{r,u}$ for $r=\frac{q+1}{4}$. The differential properties of $F_{r,u}$ are summarized in Table 1. Here, the fifth column (“Spectrum”) indicates whether the corresponding spectrum is explicitly determined (“O”) or not (“X”).
@@
q2q-2 $\pm 1$ $q\equiv 3\pmod{4}$, $p\neq 3$ $\frac{q+1}{4}$ (locally-APN) O [13]
q2q-2 $u\neq\pm 1$ $q\equiv 3\pmod{4}$, $q\neq 3$ $\chi(1+u)=\chi(u-1)=-\chi(5u\pm 3)$ 2 (APN) O [21, 20]
q2q-2 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $\chi(1+u)=\chi(u-1)\neq-\chi(5u\pm 3)$ 33 O [20]
q2q-2 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $\chi(1+u)\neq\chi(u-1)$ $\delta_{F_{r,u}}\leq 4$ O [19, 17]
22 $\pm 1$ $q\equiv 3\pmod{4}$, $\frac{q+1}{4}$ (locally-APN) X [14]
22 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $3\leq\delta_{F_{r,u}}\leq 5$ X [14]
@@
$\frac{3^{k}+1}{2}$ $\pm 1$ $p=3$, $n$ : odd, $\gcd(k,n)=1$ $\frac{q+1}{4}$ (locally-APN) X This paper
@@ 33 $\pm 1$ $q\equiv 11\pmod{12}$ $\frac{q+1}{4}$ (locally-APN) O This paper
Recently, several studies have investigated the differential and boomerang spectra of $F_{r}$ for various choices of $r$. These works consistently show that $F_{r}$ is locally-APN and has boomerang uniformity at most $2$ in all known cases. Lyu et al. [13] determined the differential and boomerang spectra of $F_{q-2}$, showing that it is locally-PN with boomerang uniformity 0 when $p=3$, and locally-APN with boomerang uniformity at most $1$ when $p\neq 3$. Mesnager and Wu [14] proved that $F_{2}$ is locally-APN with boomerang uniformity at most $2$, and further investigated its differential spectrum. They also proved that the boomerang uniformity of $F_{2}$ equals $2$, when $q\geq 9613^{2}$. However, we will show in Section 4.2 that the boomerang uniformity of $F_{2}$ equals $1$ when $p=3$. Ren and Yan [18] independently analyzed the differential spectrum of $F_{2}$ using a different approach. Koo and Kwon [10] studied $F_{r}$ for $r=\frac{q+1}{4}$, showing that $F_{r}$ is locally-PN with boomerang uniformity 0 when $q\equiv 3\pmod{8}$, and locally-APN with boomerang uniformity $2$ when $q\equiv 7\pmod{8}$. They also provided complete characterizations of the differential and boomerang spectra of $F_{r}$ in these cases. The boomerang properties of $F_{r}$ are summarized in Table 2. Here, “O” indicates that the corresponding spectrum is explicitly determined, whereas “X” indicates that it is not.
@@ 22 $q\equiv 3\pmod{4}$, large $q$ 2 X [14]
$\frac{3^{k}+1}{2}$ p=3,np=3,\;n : odd, $\gcd(k,n)=1$ $\leq 2$ X This paper
33 $q\equiv 11\;(\bmod 12)$ $\leq 2$ X This paper
$\frac{2q-1}{3}$ $q\equiv 11\;(\bmod 12)$ $\leq 2$ X This paper
@@ 22 $p=3,\;n\geq 3$ : odd 1 O This paper
In this paper, we study the differential and boomerang properties of $F_{r}$ when $q\equiv 3\pmod{4}$. In particular, we show that if the equation $(x+1)^{r}-x^{r}=b$ has at most one solution satisfying $\chi(x)=\chi(x+1)=1$ and $\gcd(r,q-1)\mid 2$, then $\delta_{F_{r}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$. This implies that $F_{r}$ is locally-APN. Furthermore, we also prove that its boomerang uniformity is at most $2$. The above condition is more likely to hold when the differential uniformity of $x^{r}$ is small, and it certainly holds when $x^{r}$ is PN. In this paper, we show that this is indeed the case for the exponents $r$ listed in Table 3 in Section 2. In addition, we further investigate the differential spectrum for $r=3$ and $r=\frac{2q-1}{3}$, and the boomerang spectrum of $F_{2}$ when $p=3$. By Lemma 2.8, the latter result can be extended to the case k=n1k=n-1 for $r=\frac{3^{k}+1}{2}$. For convenience, we recall that when $q\equiv 3\pmod{4}$, one has $\chi(-1)=-1$, and that $x^{\frac{q+1}{4}}$ provides a square root of $x\in\mathbb{F}_{q}$ whenever $\chi(x)=1$. These facts will be used throughout the paper without further mention.
@@ 2 Preliminaries
When $F(x)=x^{r}$ is a power function, multiplying $\frac{1}{a^{r}}$ on both sides of $b=F(x+a)-F(x)=(x+a)^{r}-x^{r}$, we have
@@
in this case. The notion of locally-APN functions was originally introduced in [2] for power functions in the even characteristic case. In [9], Hu et al. generalized it for odd characteristic cases, as follows.
@@
$\delta_{F}(a,b)=\delta_{F}(1,g_{a}(b))\text{ for all }b\in\mathbb{F}_{q},$ (4)
@@
Let $F$ be a function on $\mathbb{F}_{q}$ satisfying (4) for every $a\in\mathbb{F}_{q}^{*}$. Then, $F$ is called locally-APN if $\delta_{F}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}\setminus\mathbb{F}_{p}$.
@@
[6] If $q\equiv 1\pmod{4}$, then $S_{00}=\frac{q-5}{4}$ and $S_{01}=\#S_{10}=\#S_{11}=\frac{q-1}{4}$. If $q\equiv 3\pmod{4}$, then $S_{00}=\#S_{10}=\#S_{11}=\frac{q-3}{4}$ and $S_{01}=\frac{q+1}{4}$.
@@
(Theorem 5.48 of [12]) Let $f(x)=a_{2}x^{2}+a_{1}x+a_{0}\in\mathbb{F}_{q}[x]$ with $p$ odd and $a_{2}\neq 0$. Put $d=a_{1}^{2}-4a_{0}a_{2}$. Then,
$\sum_{x\in\mathbb{F}_{q}}\chi(f(x))=\begin{cases}-\chi(a_{2})&\text{ if }d\neq 0,\\ (q-1)\chi(a_{2})&\text{ if }d=0.\end{cases}$
@@
(Theorem 5.41 of [12]) Let $\eta(\cdot)$ be a multiplicative character of $\mathbb{F}_{q}$ of order $m>1$ and let $f\in\mathbb{F}_{q}[x]$ be a monic polynomial of positive degree that is not $m$-th power of a polynomial. Let $d$ be the number of distinct roots of $f$ in its splitting field over $\mathbb{F}_{q}$. Then for every $a\in\mathbb{F}_{q}$ we have
@@
In this paper, we show that for $r$ in Table 3 $F_{r}$ is locally-APN with differential uniformity $\frac{q+1}{4}$ with boomerang uniformity at most $2$. These exponents $r$ form a subclass of those satisfying $\gcd(r,q-1)=2$, for which the monomial $x^{r}$ has low differential uniformity. This is relevant to our approach, since it tends to restrict the number of solutions of $(x+1)^{r}-x^{r}=b$ for each $b\in\mathbb{F}_{q}^{*}$. Note that the differential properties summarized in Table 3 are stated for the monomial $x^{r}$, but in some cases the equivalent form $x^{r+\frac{q-1}{2}}=x^{r}\chi(x)$ is actually PN or APN. Since $F_{r+\frac{q-1}{2}}(x)=x^{r+\frac{q-1}{2}}(1+\chi(x))=x^{r}(\chi(x)+1)=F_{r}(x)$ for all $x\in\mathbb{F}_{q}$, the two forms are identical as functions.
@@ 33 $q\equiv 11\pmod{12}$ APN -
Moreover, when $r=\frac{3^{k}+1}{2}$ (the Coulter–Matthews exponent), $x^{r}$ is PN only for odd $k$ [5]. However, according to the linear equivalence established in Lemma 2.8, our results hold for all $k$.
Two functions $F$ and $F^{\prime}$ defined on $\mathbb{F}_{q}$ are said to be linearly(affine) equivalent if there exist linear(affine) permutations $L_{1}$ and $L_{2}$ over $\mathbb{F}_{q}$ such that $F^{\prime}=L_{1}\circ F\circ L_{2}$. It is well known that the differential and boomerang spectra are invariant under affine equivalence.
@@
Let $p$ be an odd prime and $n$ be odd. Then $F_{\frac{p^{k}+1}{2}}$ is linearly equivalent to $F_{\frac{p^{n-k}+1}{2}}$.
@@
Let $L(x)=x^{p^{n-k}}$. Then, for nonzero $x\in\mathbb{F}_{q}$,
$(F_{\frac{p^{k}+1}{2}}\circ L)(x)$ $=\left(x^{p^{n-k}}\right)^{\frac{p^{k}+1}{2}}(1+\chi(x^{p^{n-k}}))=x^{\frac{p^{n}+p^{n-k}}{2}}(1+\chi(x))=x^{\frac{p^{n-k}+1}{2}}\cdot x^{\frac{p^{n}-1}{2}}(1+\chi(x))$
$=x^{\frac{p^{n-k}+1}{2}}(\chi(x)+1)=F_{\frac{p^{n-k}+1}{2}}(x).$
Clearly, we can easily check that $(F_{\frac{p^{k}+1}{2}}\circ L)(0)=F_{\frac{p^{n-k}+1}{2}}(0)=0$. ∎
@@ 3 Differential Properties
In this section, we study the differential properties of $F_{r}$. We aim to count the number of solutions of
@@
Let $D_{ij}(b)$ be the number of solutions of (5) in $S_{ij}$ where $i,j\in\{0,1\}$. Then we have
$\delta_{F_{r}}(1,b)=\begin{cases}\sum D_{ij}(b)&\text{ if }b\neq 0,2,\\ \sum D_{ij}(b)+1&\text{ if }b=0\text{ or }2,\end{cases}$
where the additional term +1+1 accounts for the solutions $x=0$ and x=1x=-1, which correspond to $b=2$ and $b=0$, respectively.